OpenStudy (anonymous):

Which of the following represents the linear equation 2x = y + 3 in slope−intercept form? 2x + y = 3 2x − y = 3 y = 2x − 3 y = 2x + 3

OpenStudy (gorv):

y=mx+c 2x=y+3 2x-3=y

OpenStudy (gorv):

@christianfilms11

OpenStudy (anonymous):

Thanks.